# M.A. Basiri & S. Khanmohammadi

### v3

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. We showcase the application of our proposed method using several case studies of neuronal dynamics, where we model the macroscopic oscillation of a population of neurons using the extended Stuart-Landau (SL) equation and utilize the SINDyG method to identify the underlying nonlinear dynamics. Our extensive computational experiments validate the improved accuracy and simplicity of discovered network dynamics when compared to the original SINDy approach. Sparse Regression, Model Extraction, Network Dynamics, Neural Dynamics, Stuart-Landau Model
@@ 1 Introduction
In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. SINDyG introduces a graph-informed regularization strategy that uses prior knowledge of the network’s adjacency structure to guide model discovery. This enables more accurate and parsimonious identification of governing equations, particularly in systems where node-to-node interactions are sparse and structured. We test our method using several case studies of neuronal dynamics, where the goal is to identify nonlinear interactions between populations of neurons. Our results indicate improvements to model accuracy and simplicity when compared to the original SINDy method.
@@ 3.3 Validating the Predicted Dynamics
The coefficient mean absolute error (MAE) evaluates how closely the learned coefficients match the true coefficients. It is computed as the average absolute difference across all entries in $\bm{\xi}$. We chose MAE over mean squared error (MSE) to emphasize interpretability and robustness. MAE provides a clear indication of average deviation per term and is less influenced by outliers, which is desirable in sparse models where most terms are zero and a few incorrect activations can otherwise dominate the score. Mathematically, the coefficient MAE is defined as:
@@
The model complexity index ($\gamma$) used throughout our experiments is directly tied to the sparsity of the coefficient matrix $\mathbf{\Xi}$, as it reflects the number of nonzero terms active in the discovered dynamical equations. A sparser matrix corresponds to a lower $\gamma$, indicating a simpler and more interpretable model. Similarly, the coefficient MAE is also indirectly linked to sparsity: a higher coefficient MAE often arises when additional spurious terms are incorrectly activated in $\mathbf{\Xi}$. This observation aligns with prior findings which noted that models with poor predictive performance tend to exhibit higher complexity and reduced sparsity [14]. By reporting both $\gamma$ and coefficient MAE, we capture how closely the learned dynamics match the ground truth in both structure and accuracy.
@@ 5 Conclusion
The primary limitation of the proposed method lies in the selection and handling of appropriate basis or library functions, which could be challenging for complex systems with stochastic behavior. The size of the library functions in each point of the modeling has impact on the computational complexity of the model, which could be managed using regularization parameters. To address this, advanced optimization techniques, including sparsity-promoting algorithms and parallel computing frameworks, can be employed to reduce the computational burden without sacrificing accuracy. Additionally, exploring domain-specific knowledge to choose from library functions for specific applications, such as neuroscience, can improve computational efficiency. Lastly, while the method’s performance was validated with synthetic data, further research is needed to assess its effectiveness on real-world datasets where the true underlying dynamics are unknown. Collaborating with domain experts to acquire high-quality empirical datasets and benchmarking the method in practical scenarios could validate its robustness and utility. Nonetheless, even though we focused on the Stuart–Landau system in this study to model neuronal dynamics, the proposed method could be applied to a wide variety of nonlinear dynamical systems, such as the Kuramoto model, to understand and predict the behavior of complex systems in various fields, including neuroscience, ecology, finance, and epidemiology.

### v4

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. We tested our proposed method using several case studies of neuronal dynamics, where we modeled the macroscopic oscillation of a population of neurons using the extended Stuart-Landau (SL) equation and utilize the SINDyG method to identify the underlying nonlinear dynamics. Our extensive computational experiments validate the improved accuracy and simplicity of discovered network dynamics when compared to the original SINDy approach. The proposed graph-informed penalty can be easily integrated with other symbolic regression algorithms, enhancing model interpretability and performance by incorporating network structure into the regression process. Sparse Regression, Model Extraction, Network Dynamics, Neural Dynamics, Stuart-Landau Model
@@ 1 Introduction
In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. SINDyG introduces a graph-informed regularization strategy that uses prior knowledge of the network’s adjacency structure to guide model discovery. This enables more accurate and parsimonious identification of governing equations, particularly in systems where node-to-node interactions are sparse and structured. We tested our method using several case studies of neuronal dynamics built around the extended Stuart–Landau oscillator networks, where the goal is to identify nonlinear interactions between populations of neurons. Our results indicate improvements to model accuracy and simplicity when compared to the original SINDy method. The proposed graph-informed regularization approach is a modular framework that can be easily adapted and integrated with other symbolic regression solvers, promoting broader applicability and enhanced performance.
@@ 3.3 Validating the Predicted Dynamics
The coefficient mean absolute error (MAE) evaluates how closely the learned coefficients match the true coefficients. It is computed as the average absolute difference between the predicted coefficients and the true coefficients. We chose MAE over other metrics such as L-infinity norm to emphasize interpretability and robustness. While the L-infinity norm is useful for identifying the single worst-case deviation, it can be dominated by a single outlier and does not reflect the model’s global accuracy across all coefficients. In contrast, MAE provides a clear indication of the average deviation per term and is less influenced by outliers, which is desirable in sparse models where most terms are zero and a few incorrect activations can otherwise dominate the score. Mathematically, the coefficient MAE is defined as:
@@
The model complexity index ($\gamma$) used throughout our experiments is defined as the number of active (nonzero) terms in the coefficient matrix $\mathbf{\Xi}$. While related to sparsity, this measure is invariant to the total number of candidate terms and provides a direct and interpretable metric of model complexity. A coefficient matrix with a few active terms corresponds to a lower $\gamma$, indicating a simpler and more interpretable model. Similarly, the coefficient MAE is also indirectly linked to the number of terms: a higher coefficient MAE often arises when additional spurious terms are incorrectly activated in $\mathbf{\Xi}$. This observation aligns with prior findings which noted that models with poor predictive performance tend to exhibit higher complexity and reduced sparsity [14]. By reporting both $\gamma$ and coefficient MAE, we capture how closely the learned dynamics match the ground truth in both structure and accuracy.
@@ 5 Conclusion
The primary limitation of the proposed method lies in the selection and handling of appropriate basis or library functions, which could be challenging for complex systems with stochastic behavior. The size of the library functions in each point of the modeling has impact on the computational complexity of the model, which could be managed using regularization parameters. To address this, advanced optimization techniques, including sparsity-promoting algorithms and parallel computing frameworks, can be employed to reduce the computational burden without sacrificing accuracy. Additionally, exploring domain-specific knowledge to choose from library functions for specific applications, such as neuroscience, can improve computational efficiency. Lastly, while the method’s performance was validated using the Stuart–Landau model, further research is needed to assess the effectiveness of the proposed approach on other nonlinear dynamical systems, such as the Kuramoto and FitzHugh–Nagumo models. Nonetheless, the proposed graph-informed penalty term is very general and could be integrated with other symbolic regression algorithms to better understand and predict the behavior of complex systems across various domains, including neuroscience, ecology, finance, and epidemiology.